Points D, E, and F are collinear with E between D and F. DE=15, EF=x+17, and DF=3x-10. Find EF and DF.
Answer by richwmiller(17219)

DE+EF= DF
15+x+17=3x-10
32+x=3x-10
42=2x
21=x
EF=x+17
EF=21+17=38
DF=3x-10
DF=3*21-10
DF=63-10=53
DE+EF= DF
check
15+38=53
ok
53=53